OpenStudy (anonymous):

log_5 (2x-5) =-4 log_5 (2x-5) =-4 @Mathematics

OpenStudy (anonymous):

exponentiate both sides by 5to cancel out the log\[5^{\log_{5}(2x-5)}=5^{-4} \] \[2x-5=1/625\]

OpenStudy (anonymous):

\[2x-5+5=1/625+5\]\[2x=3126/625\]\[x=1563/625\]